OpenStudy (anonymous):

The number of passengers using a railway fell from to during a -year period. Find the annual percentage decrease over this period.

OpenStudy (anonymous):

196707 to 185184 over 5 years

OpenStudy (anonymous):

196707=100% 185184=x% 196707/100=185184/x 196707x=18518400 x=94.14% 100%-94.14%=5.86% 5.86%/5years=1.17% It went down approximately 1.17% each year

OpenStudy (anonymous):

ty :D

OpenStudy (anonymous):

np :)